There are 10 girls and 7 boys in a class and need to select 4 students as a school representative. In how many ways can we select of 2 girls and 2 boys?


1
Expert's answer
2022-06-02T06:22:06-0400

We can select 2 out of 7 boys in "\frac{7!}{2!5!}= 21" ways.

We can select 2 out of 10 girls in "\frac{10!}{2!8!}= 45" ways.


We can select 2 boys and 2 girls in "21\cdot 45=945" ways.


Answer: 945 ways.